Find the general solution of the lagrange's equation 2yzp+zxq=3xy?


1
Expert's answer
2022-02-14T17:42:00-0500

Solution:

"2yzp+zxq=3xy"

"\frac{dx}{2yz}=\frac{dy}{zx}=\frac{dz}{3xy}"

"3ydy=zdz"

"3y^2-z^2=c_1"

"xdx=2ydy"

"x^2-2y^2=c_2"

"F(c_1,c_2)=F(3y^2-z^2,x^2-2y^2)=0"